Assume a company has a current stock price of 80 and will pay 2.8 as dividend in one year; its equity cost of capital is 14%. what price must you expect the stock to sell for immediately after the firm pays the dividend in one year to justify its current price? (Round your answer to two decimal digits) 


A. 88,40 

B. 82,80 

C. 91,20 

D. 94,00 

E. 66,00 

F. None of the above is correct.
Answer #1

Current stock price = 80

Cost of capital = 14% = 0.14

Value of this price at the end of year 1 = 80 * (1+0.14) = 80*1.14 = $91.20

Dividend to be paid at the end of year 1 = $2.80

Value of stock after payment of dividend = $91.20 + $2.80 = $94

Hence, correct answer is D. 94